OpenStudy (anonymous):

find the x-intercepts x^2+7x+36

OpenStudy (anonymous):

you use the quadratic formula \[-b \pm \frac{ \sqrt{b^2 -4ac} }{2a}\] Once you plug it in you'll get x=\[\frac{ -7\pm i \sqrt{95} }{ 2 }\]